# Rethinking Broken Object Level Authorization Attacks Under Zero Trust Principle

### v1

@@ 1. Introduction
BOLARAY (CCS’24, Oct) (Huang et al., 2024) is the SOTA method based on authorization model to detect BOLA attacks. By analyzing actual BOLA vulnerabilities in open-source applications, the BOLARAY team summarized the four most common object-level authorization models. However, artificial inductive authorization models are not suitable for all application systems. As shown in Fig.1, BOLARAY determines the Article table as the ownership model, in which an article can only be updated/deleted by its owner. The authorization models summarized by BOLARAY cannot support the BOLA vulnerability detection of Select statements, and are also invalid in some scenarios. for example, API5 (user likes article) in Fig.1, update statement can add likes to all articles, but BOLARAY still determines that API5 can only update the user’s own article. The artificial inductive authorization model does not fully conform to the data flow of resource ID propagation in web applications (akamai, 2024). Authorization models that are out of context semantics will reduce the accuracy of attack detection.
We rethink the nature of BOLA attacks- horizontal privilege escalation. The defense problem of the BOLA attack is transformed into isolating the injection point (resource ID) to the minimum permission range that conforms to the system logic. The resource ID parameter serves as both attack surfaces (Sengupta et al., 2020)—vulnerable to exploitation through horizontal privilege escalation—and protection surfaces (Campbell, 2020), which should enforce a minimum permission boundary for each user. Effective defense against BOLA attacks requires determining the proper authorization interval for resource IDs, ensuring users operate within the minimum necessary permissions (Jero et al., 2021).
Zero Trust (Syed et al., 2022), grounded in the principle of “never trust, always verify” (Samaniego and Deters, 2018), emphasizes fine-grained, identity-based access control to mitigate risks associated with unauthorized lateral movement. Micro-segmentation (MSG), a crucial component of zero trust, creates a security closed-loop by considering multiple factors—such as user roles, processes, and access contexts—to define the minimal security boundary for resource IDs based on API context data flow. This not only aligns with the API’s authorization logic but also enforces the principle of minimum permissions, making MSG an effective strategy for defending against BOLA attacks.
@@
To the best of our knowledge, this is the first work to apply zero trust principles to the discovery of attack surfaces and the refinement of data-level permissions in RESTful APIs. Among the 94 RESTful APIs analyzed in the selected GitHub projects, BolaZ achieved recall rates of 97% for attack injection points and 87% for authorization intervals, with only one false positive. The experimental results are promising; BolaZ successfully defended against known BOLA vulnerabilities in our benchmark and identified 35 new BOLA vulnerabilities in 10 real-world projects.
@@ 2.3. Micro-Segmentation (MSG)
Figure 2. Attack and defense
@@ 3.2. Terminology
Figure 4. API types
@@ 3.4.1. Resource storage based on database
There are many forms of resource storage in RESTful APIs, such as databases, Hadoop, and Spark. The database is the main carrier for resource storage. The current SOTA method-BOLARAY only supports database storage. Therefore, BolaZ considers the database as a resource storage carrier for BOLA attack detection.
@@ 4.2. P-API Identification and MSG Interval Generation
The P-API is the upstream of the resource ID propagation, and generates the resource ID’s MSG interval. Therefore, BolaZ’s first step is to identify the P-API and get the MSG interval of the resource ID.
BolaZ argues that Select statments represent MSG intervals. As shown in Fig.6, the three Select statements produce MSG intervals for the resource ID of blog (blog_id), respectively. The MSG interval a contains the blog_id created by user A himself. The MSG interval b contains the blog_id created by user b himself. The MSG interval c contains all blog_id. The false MSG interval only reads the amount of data in the blog table and does not generate resource ID.
@@ 4.2.1. P-API Identification
As shown in Fig.8a, BOLAZ tracks the propagation of tokens in the API layer and SQL layer, and obtains the Select statements associated with the API return value. By analyzing the Abstract syntax tree (AST) of the Select statment, BolaZ determines that the Select statement obtains the primary key id of the blog table. Only when the associated Select statement obtains the resource ID columns of the table, BolaZ determines the API as P-API.
Figure 7. FP-API Identification. SQLa produces the same blog id as the API parameters, it does not create a new blog id. BolaZ considers the API for creating and consuming the same resource ID is not a P-API, but a FP-API.
@@ 4.2.2. MSG Interval Generation
Integrity of the MSG interval. Select statements represents MSG intervals. Still, there may be some conditions in Select statements to reduce the range of MSG intervals. As shown in Listing 1(1), the SQL statement obtains the user’s address. “Detail” condition corresponds to the keyword search function of the address, and the user can make the detail condition cover the entire table without entering the keyword. “Limit” corresponds to the address paging function, and users can also obtain all data through the page turning function. Conditions of this type will reduce MSG intervals. User_id is the foreign key of the addr table, which is a necessary condition to determine MSG intevals. Therefore, BolaZ preserves only the foreign key and non-user input conditions in Select statements, as shown in Listing 1(2).
@@
Third, there may be a parent-child dependency between resource IDs in P-APIs. Obtaining the MSG interval of the parent resource ID is the premise of calculating the MSG interval of the sub-resource ID. As shown in Fig.10, the API uses SQLa and SQLb to produce MSG intervals of blog ID and discuss ID, respectively. SQLa provides blog IDs for SQLb, and SQLb depends on SQLa. BolaZ tracks the dependency of resource IDs within P-API between SQL statements.
@@ 4.3. C-API Identification and BOLA Attack Injection Point Discovery
BOLA attack injection points exist in user-controlled resource ID parameters, which eventually propagate to the SQL statement’s primary or foreign key. Since the resource ID parameter does not change in C-API propagation process (detailed in  3.4.3), BolaZ also uses taint tracking technology (Kim et al., 2014) to analyze which API parameters are propagated to SQL statements, and determines which resource ID is associated with API parameters by analyzing the AST of SQL statements.
As show in Fig.8b, BolaZ uses API parameters (blogId) as taint sources and SQL statements as taint propagation endpoints. BolaZ found that blogId eventually spread to the id condition of Delete statement to delete blog. BolaZ determines that the deleteBlog{blogId} is C-API, and the blogId parameter consumes the primary key id of the blog table.
@@ 4.4. P-API and C-API Data Flow Association
As show in Fig.8c. When the page is loaded, the Javascript code calls the P-API (/getUserBlog) to get the blog created by the user himself. The HTML code renders the blog resources returned by the P-API and provides a click event so that the user can call the deleteBlog function. When the user initiates a click event in the page, the Javascript code calls the C-API (/deleteBlog/{blogId}) to delete the corresponding blog resource. BolaZ takes P-API as the starting point of taint tracking and C-API as the end point of taint tracking, determines which P-API provides resource IDs for the C-API.
@@ 4.4.1. Identify The Data Flow of P-APIs And C-APIs
BolaZ captures the propagation of resource ID within and between HTML pages. First, we analyze the propagation of resource ID within the HTML page. As shown in Fig.11, BolaZ divides the internal propagation of the page into four data flow patterns, with P-API as the starting point of data flow propagation. C-API will terminate the data flow after consuming the resource ID. The router is used to realize the function of jumping from one page to another page, and the parameters can be carried during the jumping process.
@@ 4.4.2. Generation of MSG Interval Set
The foreign key is inserted into the database by the C-API of the resource creation type (POST), so the MSG range of the foreign key is created by the P-API that provides the resource ID for the POST C-API. As shown in Fig.12, in data stream 1, the P-API of Resource B passes the resource ID of the primary key to the POST C-API of Resource A, and the C-API stores the primary key of Resouce B as the foreign key of Resource A. In data stream 2, the P-API propagates the resource ID of the foreign key to the C-API. The MSG interval of the foreign key resource ID in data stream 2 comes from the P-API in data stream 1. Therefore, BolaZ converts the MSG interval of the foreign key into the MSG interval of the primary key of the P-API associated with the POST C-API through data stream backtracking, as shown in data stream 3.
@@ 4.4.3. Optimization of MSG Interval Set
There may be subset or intersection relationships between the internal elements of the MSG interval set, and performing repeated interval queries will result in considerable performance cost. Therefore, BolaZ defines the merging rules of Select statements to reduce its number under the premise of ensuring that the MSG interval does not change, thus reducing the number of Select statements and the query of repeated intervals.
@@ 5. Implementation
In this chapter, we detail the implementation of BolaZ. As shown in Fig.14, BolaZ includes offline analysis and runtime authorization. While numerous frameworks like SpringBoot, Flask, and Django are available for implementing RESTful APIs, BolaZ’s implementation does not currently support all of them. After reviewing these popular frameworks on GitHub, we found that SpringBoot had the highest usage and star ratings. We also searched for the frequency of use of frameworks in the industry (TheirStack, 2025), with companies using SpringBoot surpassing frameworks such as Diango, Rails, and ThinkPHP. Besides, SpringBoot integrates features such as authentication and authorization to ensure the security of resources and has become the most popular framework in the Java world (masteringbackend, 2024). As a result, BolaZ was implemented using SpringBoot-based applications to demonstrate its capabilities, leveraging the framework’s widespread adoption and strong support for RESTful services.
Regardless of the framework or programming language, the propagation behavior of resource IDs remains consistent. BolaZ needs to determine how the framework receives the parameters, how the resource ID is passed, and how it is returned. These features are consistent in other frameworks. We do not rely on SpringBoot’s unique features in implementing BolaZ. Therefore, BolaZ ’s approach can be adapted for use with other frameworks.
@@ 5.2. Runtime Authorization
BolaZ checks the authorization of resource ID according to MSG intervals generated by offline analysis when the system is running. Resource IDs used by users must exist in MSG intervals, otherwise, the C-API call will be blocked. We consider that when the database data size is large, if database queries are performed at each C-API authorization check, it will lead to high latency. C-APIs will be called in a short time after users call P-APIs. Hence, MSG engine caches resource IDs produced by P-APIs. If the cache is not hit, a database query is performed. A cache or database hit means the C-API has access to the resource ID.
6. Experiment
@@ 6.1. Research Questions
BOLARAY is the most closely relevant work to BolaZ and is the SOTA method for detecting BOLA vulnerabilities. Therefore, we compared the vulnerability detection effect of BOLARAY and BolaZ in real world projects.
@@ 6.2.1. How to collect projects
Since BolaZ needs to analyze the source code, we searched for repositories that depend on SpringBoot on GitHub (spring-projects, 2025), SourceCodeExamples (sourcecodeexamples, 2025) and LibHunt (libhunt, 2025) and filtered them by application types, authentication modes, and logic control modes.
We selected 10 open-source projects from nearly 100 real-world projects. Not all of the nearly 100 projects meet the Restful specification, and BolaZ does not support the analysis of this type of API in the implementation process. So we will remove these items in the project select process.
The 10 projects in the experiment cover various application types, two authentication modes, and two logic control modes. 1). Application type. There are similarities in the business types of open-source projects. The project we selected contains Blog, Bookstore, E-commerce, Content Management System, Online Examination System, People Information System etc. We screen systems with different business logic so that BolaZ can cover a variety of business scenarios. 2). User authentication mode. User authentication modes include token and cookie, so the system we have chosen covers both user authentication modes. 3). Logical control modes (detailed in  3.4.2), as shown in Table 3. Based on these three, the selected projects are sufficient and can effectively cover our scope of problems. We focus on the API data flow, which requires a good coverage of business types. It is of little significance to involve more projects.
@@ 6.2.2. RQ1 & RQ2
There is no tool to achieve the classification and association of API. To construct the dataset’s ground truth, we need deployed projects locally, and manually counted and identified P-API, C-API and the relationship. The whole process is complex and time-consuming. So we selected five systems from Table 3 based on application types, logical control mode, and user authentication modess to contribute a benchmark to verify the effectiveness of BolaZ, as shown in Table 4. Sufficient to verify the effectiveness of BolaZ.
We manually analyzed the API categories and relationships to build the ground truth data. First, we deployed and ran the project locally, logging in as a regular user to access all functionalities. Simultaneously, we used Fiddler (telerik, 2024) to capture the API list. Next, we examined the API server’s source code. Given that the current architecture follows a three-tier model, we were able to quickly categorize the APIs. We then analyzed the front-end source code to trace the propagation paths of resource IDs both within and across pages. Our team, divided into two groups, analyzing the project’s source code independently, followed by cross-validation of the results. Across the five projects, we identified a total of 40 P-APIs and 54 C-APIs.
@@ 6.2.3. RQ3-1
We searched the CVE vulnerability database for BOLA (a.k.a., IDOR) vulnerabilities. The results indicate that only a few from Java projects (e.g., SpringBoot). Finally, we screened out three vulnerabilities corresponding to the three modes of BOLA attack from the CVE-2023-36100 (CVE, 2023b) and CVE-2023-32310 (CVE, 2023a).
@@ 6.4.1. Setup
Table 5. Effectiveness of API Classification and Association, P denote P-API, C denote C-API, R denote Relation.
@@ 6.4.2. Results
The experimental results of BolaZ are reported in Table 5. BolaZ achieved recall rates of 97% for API classification and 87% for API association, with one false positives. The precision of API classification and association is high because the resource ID, as the unique identification of the resource, is rarely modified in the process of data flow propagation. BolaZ can categorize and associate APIs very accurately without interrupting the taint tracking. We also analyze the reasons for the failure of API classification. By examining the log, we found that CodeQL generated data flow interruption at the Arrays.asList method during the taint tracking process, failing P-API identification of the Newbellmall project. The Blog project also failed C-API recognition due to data flow interruption.
@@
BolaZ generated the only false positive in analyzing API relationships in the online-exam project. After analyzing the source code, we found that the reason was that the client code modified the resource ID generated by P-API and passed it to the downstream C-API, resulting in BolaZ incorrectly associating the context of P-API and C-API. BOLAZ needs to analyze the source code to obtain the modification logic of the resource ID to solve this type of false positives. The problem is transformed into extracting operational semantics from source code.

### v2

@@ 1. Introduction
BOLARAY (CCS’24, Oct) (Huang et al., 2024) is the state-of-the-art (SOTA) method for detecting BOLA attacks. By analyzing real-world BOLA vulnerabilities in open-source applications, BOLARAY identifies four common object-level authorization models. However, these models, being manually summarized, lack adaptability to the diverse and evolving nature of modern web applications. As illustrated in Fig.1, BOLARAY classifies the Article table under the ownership model, where an article can only be updated or deleted by its owner. This model fails to support SELECT operations. For instance, API5 in Fig.1 involves an UPDATE statement that allows users to add likes to articles. However, BOLARAY incorrectly assumes that articles can only be updated by their owners. This rigid interpretation overlooks the actual propagation patterns of resource identifiers in web applications (akamai, 2024), leading to semantic mismatches. As a result, such out-of-context authorization logic limits BOLARAY’s ability to comprehensively detect BOLA vulnerabilities.
We revisit the essence of BOLA attacks—horizontal privilege escalation and reconceptualize their defense as a problem of minimally scoped authorization. Specifically, defending against BOLA attacks involves isolating the injection point (i.e., the resource ID) within the narrowest permission boundary that aligns with the application’s logic. The resource ID parameter functions dually as an attack surface vulnerable to horizontal privilege escalation (Sengupta et al., 2020), and as a protection surface responsible for enforcing least privilege access control (Campbell, 2020). Effective mitigation requires accurately determining the authorization interval for each resource ID, ensuring users can only access or manipulate resources within their minimum necessary permissions (Jero et al., 2021).
Zero Trust, grounded in the principle of “never trust, always verify” (Samaniego and Deters, 2018), emphasizes fine-grained, identity-based access control to mitigate risks associated with unauthorized lateral movement. Micro-segmentation (MSG), a crucial component of zero trust, creates a security closed-loop by considering multiple factors—such as user roles, processes, and access contexts—to define the minimal security boundary for resource IDs based on API context data flow. This not only aligns with the API’s authorization logic but also enforces the principle of minimum permissions, making MSG an effective strategy for defending against BOLA attacks.
@@
To the best of our knowledge, this is the first work to apply Micro-segmentation (MSG), which is derived from zero trust principles, to the discovery of attack surfaces and the refinement of data-level permissions in RESTful APIs. Among the 94 RESTful APIs analyzed in the selected GitHub projects, BolaZ achieved recall rates of 97% for attack injection points and 87% for authorization intervals, with only one false positive. The experimental results are promising; BolaZ successfully defended against known BOLA vulnerabilities in our benchmark and identified 35 new BOLA vulnerabilities in 10 real-world projects.
@@
The rest of this paper is organized as follows. Section 2 introduces background knowledge on RESTful APIs, BOLA attacks and Micro-Segmentation (MSG). Section 3 demonstrates how the problem of BOLA vulnerability detection is formulated at the methodological level. In Section 4, the technical details of the proposed novel approach, BolaZ, are described. Section 5 shows the details of the implementation. Section 6 describes the experiments and discusses the results. Section 7 provides discussion and future work. Relevant literature is summarized in Section 8. Section 9 concludes this work.
@@ 2.3. Micro-Segmentation (MSG)
Figure 2. A BOLA attack on an order details API and its prevention by Micro-Segmentation (MSG)
@@ 3.2. Terminology
Figure 4. Classification of APIs based on rid production and consumption
@@ 3.4.1. Resource Storage Based on Database
There are many forms of resource storage in RESTful APIs, such as databases, Hadoop, and Spark. The database is the main carrier for resource storage. Therefore, BolaZ considers the database as a resource storage carrier for BOLA attack detection, the same as BOLARAY (Huang et al., 2024).
@@ 4.2. P-API Identification and MSG Interval Generation
The P-API is the upstream of the resource ID propagation and generates the resource ID’s MSG interval. Therefore, BolaZ’s first step is to identify the P-API and get the MSG interval of the resource ID.
BolaZ argues that SELECT statements represent MSG intervals. As shown in Fig.6, the three SELECT statements produce MSG intervals for the resource ID(blog_id) of blog, respectively. The MSG interval a contains the blog_id created by user A himself. The MSG interval b contains the blog_id created by user b himself. The MSG interval c contains all blog_id. The false MSG interval only reads the amount of data in the blog table and does not generate resource ID.
@@ 4.2.1. P-API Identification
As shown in Fig.8a, BOLAZ tracks the propagation of tokens in the API layer and SQL layer, and obtains the SELECT statements associated with the API return value. By analyzing the Abstract syntax tree (AST) of the SELECT statement, BolaZ determines that the SELECT statement obtains the primary key id of the blog table. Only when the associated SELECT statement obtains the resource ID columns of the table, BolaZ determines the API as P-API.
Figure 7. FP-API Identification. SQL-(a) produces the same blog ID as the API parameter, not creating a new blog ID. BolaZ considers the API for creating and consuming the same resource ID as a P-API, not an FP-API.
@@ 4.2.2. MSG Interval Generation
Integrity of the MSG interval. SELECT statements represent MSG intervals. Still, there may be some conditions in SELECT statements to reduce the range of MSG intervals. As shown in Listing 1(1), the SQL statement obtains the user’s address. “Detail” condition corresponds to the keyword search function of the address, and the user can make the detail condition cover the entire table without entering the keyword. “Limit” corresponds to the address paging function, and users can also obtain all data through the page turning function. Conditions of this type will reduce MSG intervals. User_id is the foreign key of the addr table, which is a necessary condition to determine MSG intervals. Therefore, BolaZ preserves only the foreign key and non-user input conditions in SELECT statements, as shown in Listing 1(2).
@@
Third, there may be a parent-child dependency between resource IDs in P-APIs. Obtaining the MSG interval of the parent resource ID is the premise of calculating the MSG interval of the sub-resource ID. As shown in Fig.10, the API uses SQL-(a) and SQL-(b) to produce MSG intervals of blog ID and discuss ID, respectively. SQL-(a) for SQL-(b), and SQL-(b) depends on SQL-(a). BolaZ tracks the dependency of resource IDs within P-API between SQL statements.
@@ 4.3. C-API Identification and BOLA Attack Injection Point Discovery
BOLA attack injection points exist in user-controlled resource ID parameters, which eventually propagate to the SQL statement’s primary or foreign key. Since the resource ID parameter does not change in the C-API propagation process (detailed in  3.4.3), BolaZ also uses taint tracking technology (Kim et al., 2014) to analyze which API parameters are propagated to SQL statements, and determines which resource ID is associated with API parameters by analyzing the AST of SQL statements.
As shown in Fig.8b, BolaZ uses API parameters (blogId) as taint sources and SQL statements as taint propagation endpoints. BolaZ found that blogId eventually spread to the id condition of Delete statement to delete blog. BolaZ determines that the deleteBlog{blogId} is C-API, and the blogId parameter consumes the primary key id of the blog table.
@@ 4.4. P-API and C-API Data Flow Association
As shown in Fig.8c. When the page is loaded, the Javascript code calls the P-API (/getUserBlog) to get the blog created by the user himself. The HTML code renders the blog resources returned by the P-API and provides a click event so that the user can call the deleteBlog function. When the user initiates a click event in the page, the Javascript code calls the C-API (/deleteBlog/{blogId}) to delete the corresponding blog resource. BolaZ takes P-API as the starting point of taint tracking and C-API as the end point of taint tracking determines which P-API provides resource IDs for the C-API.
@@ 4.4.1. Identify The Data Flow of P-APIs And C-APIs
BolaZ captures the propagation of resource ID within and between HTML pages. First, we analyze the propagation of resource ID within the HTML page. As shown in Fig.11, BolaZ divides the internal propagation of the page into four data flow patterns, with P-API as the starting point of data flow propagation. C-API will terminate the data flow after consuming the resource ID. The router is used to realize the function of jumping from one page to another, and the parameters can be carried out during the jumping process.
@@ 4.4.2. Generation of MSG Interval Set
The foreign key is inserted into the database by the C-API of the resource creation type (POST), so the MSG range of the foreign key is created by the P-API that provides the resource ID for the POST C-API. As shown in Fig.12, in data stream 1, the P-API of Resource B passes the resource ID of the primary key to the POST C-API of Resource A, and the C-API stores the primary key of Resource B as the foreign key of Resource A. In data stream 2, the P-API propagates the resource ID of the foreign key to the C-API. The MSG interval of the foreign key resource ID in data stream 2 comes from the P-API in data stream 1. Therefore, BolaZ converts the MSG interval of the foreign key into the MSG interval of the primary key of the P-API associated with the POST C-API through data stream backtracking, as shown in data stream 3.
@@ 4.4.3. Optimization of MSG Interval Set
There may be subset or intersection relationships between the internal elements of the MSG interval set, and performing repeated interval queries will result in considerable performance costs. Therefore, BolaZ defines the merging rules of SELECT statements to reduce its number under the premise of ensuring that the MSG interval does not change, thus reducing the number of SELECT statements and the query of repeated intervals.
@@ 5. Implementation
In this section, we detail the implementation of BolaZ, a tool for detecting BOLA vulnerabilities in SpringBoot-based web applications. As shown in Fig.14, BolaZ includes offline analysis and runtime authorization.
@@ 5.2. Runtime Authorization
BolaZ checks the authorization of resource ID according to MSG intervals generated by offline analysis when the system is running. Resource IDs used by users must exist in MSG intervals, otherwise, the C-API call will be blocked. We consider that when the database data size is large, if database queries are performed at each C-API authorization check, it will lead to high latency. C-APIs will be called in a short time after users call P-APIs. Hence, the MSG engine caches resource IDs produced by P-APIs. If the cache is not hit, a database query is performed. A cache or database hit means the C-API has access to the resource ID.
6. Evaluation
@@ 6.1. Research Questions
BOLARAY is the most closely relevant work to BolaZ and is the SOTA method for detecting BOLA vulnerabilities. Therefore, we compared the vulnerability detection effect of BOLARAY and BolaZ in real-world projects.
@@ 6.2.1. How to collect projects
As the implementation of BolaZ is on the SpringBoot framework in Java language, we searched for repositories that depend on SpringBoot on GitHub (spring-projects, 2025), SourceCodeExamples (sourcecodeexamples, 2025) and LibHunt (libhunt, 2025) and filtered them by application types, authentication modes, and logic control modes.
The selected 10 projects cover various application types, two authentication modes, and two logic control modes. 1). Application type. There are similarities in the business types of open-source projects. The project we selected contains Blog, Bookstore, E-commerce, Content Management System, Online Examination System, People Information System, etc. We screen systems with different business logic so that BolaZ can cover a variety of business scenarios. 2). User authentication mode. User authentication modes include tokens and cookies, so the system we have chosen covers both user authentication modes. 3). Logical control modes (detailed in  3.4.2), as shown in Table 3. Based on these three, the selected projects are sufficient and can effectively cover our scope of problems. We focus on the API data flow, which requires a good coverage of business types. It is of little significance to involve more projects.
@@ 6.2.2. RQ1 & RQ2
There is no tool to achieve the classification and association of API. To construct the dataset’s ground truth, we need to deploy projects locally and manually count and identify P-API, C-API and the relationship. The whole process is complex and time-consuming. So we selected five systems from Table 3 based on application types, logical control mode, and user authentication modes to contribute a benchmark to verify the effectiveness of BolaZ, as shown in Table 4. Sufficient to verify the effectiveness of BolaZ.
We manually analyzed the API categories and relationships to build the ground truth data. First, we deployed and ran the project locally, logging in as a regular user to access all functionalities. Simultaneously, we used Fiddler (telerik, 2024) to capture the API list. Next, we examined the API server’s source code. Given that the current architecture follows a three-tier model, we were able to quickly categorize the APIs. We then analyzed the front-end source code to trace the propagation paths of resource IDs both within and across pages. Our team was divided into two groups, analyzing the project’s source code independently, followed by cross-validation of the results. Across the five projects, we identified a total of 40 P-APIs and 54 C-APIs.
@@ 6.2.3. RQ3-1
We searched the CVE vulnerability database for BOLA (a.k.a., IDOR) vulnerabilities. The results indicate that only a few are from Java projects (e.g., SpringBoot). Finally, we screened out three vulnerabilities corresponding to the three modes of BOLA attack from the CVE-2023-36100 (CVE, 2023b) and CVE-2023-32310 (CVE, 2023a).
@@ 6.4.1. Setup
Table 5. Effectiveness of API Classification and Association, P denotes P-API, C denotes C-API, R denotes Relation.
@@ 6.4.2. Results
The experimental results of BolaZ are reported in Table 5. BolaZ achieved recall rates of 97% for API classification and 87% for API association, with one false positive. The precision of API classification and association is high because the resource ID, as the unique identification of the resource, is rarely modified in the process of data flow propagation. BolaZ can categorize and associate APIs very accurately without interrupting the taint tracking. We also analyze the reasons for the failure of API classification. By examining the log, we found that CodeQL generated data flow interruption at the Arrays.asList method during the taint tracking process, failing P-API identification of the Newbellmall project. The Blog project also failed C-API recognition due to data flow interruption.
@@
BolaZ generated the only false positive in analyzing API relationships in the online-exam project. After analyzing the source code, we found that the reason was that the client code modified the resource ID generated by P-API and passed it to the downstream C-API, resulting in BolaZ incorrectly associating the context of P-API and C-API. BOLAZ needs to analyze the source code to obtain the modification logic of the resource ID to solve this type of false positive. The problem is transformed into extracting operational semantics from the source code.